Suppose that before any tax on labor income, the equilibrium wage in a market is $10 per hour.
Once the government imposes a $3 per hour tax on labor income, the equilibrium wage rises to $11 per hour. What does that reveal about the elasticities of supply and demand for labor?
      
A) Demand is more elastic than supply.
      
B) Both the supply and demand are inelastic.
      
C) Supply is more elastic than demand.
      
D) Supply and demand are equally elastic.
      
E) Nothing because more information is needed to learn about the elasticities.
Answer #1

After tax income of worker=11-3=$8

Burden of tax passed to workers=10-8=$2

Burden of tax passed to demanders=11-10=$1

We can see that amount of tax burden passed in more in case of labor (supply side)

It means that demand is more elastic in this case.

Correct option is

A) demand is more elastic than supply.